OpenStudy (sam_pi):

I need help with solving these two 2 problems Given that log_a (2) = 0.301, log_ a (5) = 0.699, and log _a ( 6 ) = 0.778 find log_a(4) and log_a(1/6) I don’t understand how you are suppose to find the answer using the change of base formula

OpenStudy (sam_pi):

do you have any explanations or suggestions on how to begin to solve them?

OpenStudy (fourier):

\[\log_a 4 = \log_a 2^2 = 2\times \log_a 2 = ?\]

OpenStudy (fourier):

I have used the property \[\log_a x^n = n\times \log_a x\]